OpenStudy (anonymous):

The table below shows the ticket rates for whale watching trips offered by Ocean Tours. Age Ticket Price Under three years free Three to twelve years $36 Over twelve years $48 On a certain day, the company took 92 people on whale watching trips. There were 40 children aged twelve and under, of which some children were under three years. If x represents the number of children undr three years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day? @magbak

OpenStudy (anonymous):

(92 - x)36 + (92 – 40) ∙ 48 = C (40 + x)36 + (92 – (40 + x)) ∙ 48 = C (92 + x)36 + (92 – (40 + x)) ∙ 48 = C (40 – x)36 + (92 – 40) ∙ 48 = C

OpenStudy (magbak):

I am here.

OpenStudy (anonymous):

@magbak help

OpenStudy (magbak):

PLEASE DO NOT RUSH ME OK!!!!

OpenStudy (anonymous):

okay sorry @magbak

OpenStudy (magbak):

I am going to need help on this can I get some assistance or not.

OpenStudy (anonymous):

@GoldPhenoix

OpenStudy (magbak):

NO I HAVE MORE SOFISTICATED PEOPLE LEAVE THIS TO ME OK.

OpenStudy (anonymous):

okay

OpenStudy (magbak):

@jim_thompson5910 @Hero @Loser66 @satellite73 @whpalmer4

OpenStudy (goldphenoix):

40 - x represent the amount of kids under 3. Does that make sense so far?

OpenStudy (anonymous):

kinda @GoldPhenoix

OpenStudy (anonymous):

nvm ill just guess

OpenStudy (goldphenoix):

Actualy. That represent the amount of kids over 3*

OpenStudy (goldphenoix):

What is it then?

OpenStudy (anonymous):

the third one

OpenStudy (goldphenoix):

It's not the third one.

OpenStudy (anonymous):

is it one of the 92 or the 40

OpenStudy (goldphenoix):

What do you mean?

OpenStudy (anonymous):

like cant you cross out 2

OpenStudy (loser66):

I vote for the last one

OpenStudy (goldphenoix):

It is the last one. >_> Think about it. (92-40) represent the amount of student over 12. Since 40 represent 12 years old and less. Now you multiply that by 48 to get the cost. Make sense?